# MRNaB: Mixed Reality-based Robot Navigation Interface using Optical-see-through MR-beacon

### v1

@@ V Experiment
To evaluate the effectiveness of our system, we conducted an experiment to compare between our proposed system with the 2D system by using the computer display and mouse as the baseline as it is more used in daily life. The same experiment would be conducted twice, each for 2D system and our system (MRNaB). Fourteen participants were involved in this experiment, comprising 7 women and 7 men with their ages ranging from 20 to 28. Most of our participants didn’t have VR/AR/MR nor ROS experience before. The order of the experiments for the conventional method and the proposed method was randomized.
@@ V-A Experiment Setup
Fig. 9 shows the experiment environment. We made the environment to be similar to the household setting. For the task, we asked the participant to navigate the robot using beacon to a certain area, where the allowed area and the direction were decided beforehand. This area was further divided into 4 stages, which were stage 1, stage 2, stage 3, and stage 4.
In stage 1, we wanted to cover a reasonably spacious area, that resembled a kid’s area where there is usually no hint from the map created by SLAM. In stage 2, we reduced the size of the area, however, we gave the opportunity for the participant to be able to see the hint reflected from the map created by SLAM. This stage resembled the coffee station in the household environment. In stage 3, we kept the similar size of the area, while starting to reduce the hint reflected from the map created by SLAM. This stage resembled the working seat. Lastly in stage 4, we removed any hint from the area surrounding the target area, thus making the participant needed to use the information from the previous stage to create the MR-beacon in 2D system. This stage resembles the case where the robot needs to get the water from the water leakage from the roof.
Fig. 10 shows the experiment flow. Each of the experiment was conducted as follows: First, participants would be given an explanatory video on how to use each of the systems. After that, we helped the participants to get used to each of the main functions. Next, we let the participants to try the system themselves. Last was the real experiment. Every time the robot was navigated outside the area, participants would be required to move the beacon and navigate it again.
@@ V-B Evaluation Indices
TABLE II: Effectiveness Comparison by Action Number
@@
Evaluation indices would be divided into 2, which were objective indices and subjective indices. For objective indices, we would first measure the total action number before navigation measurement. The total action number refers to the total number of clicks on the Add and Move buttons. This was the total action needed before participants finally be able to navigate the robot to the correct place. Second, we would measure navigation number measurement results per task, this was the total navigation done by user per stage.
For subjective indices, we would measure usability by system usability scale (SUS) questionnaire [35], this was a method in which 10 questions regarding system usability would be given and then be rated on a scale of 0 to 4, with the sum multiplied by 2.5 and rated on a scale of 0 to 100.
V-C Experiment Result
@@
TABLE III: Effectiveness Comparison by Navigation Number
@@
V-C1 Total Action Number Before Navigation Measurement Result
The result is shown in Fig. 11. Given that the $p$-values for stages 1, 2, 3, and Overall were below 0.05, we concluded there was statistical significance across these stages. Referencing Table II, it was evident that participants required less beacon actions to navigate the robot accurately using our system compared to a traditional 2D system where our system only required 1.59 tries on average while the 2D system required 2.95 tries on average. We posit that these changes came because, in our system, participants were able to navigate the robot directly in the real world compared to the 2D system thus making it easier for participants to navigate the robot directly to the desired location.
@@ V-C2 Navigation Number Measurement Results Per Task
The result is shown in Fig. 12. Given that the $p$-values for all stages were lower than 0.05, we concluded that there is statistical significance in the data presented. Referencing Table II, it was evident that participants required less navigation for the robot to reach its destination using our system compared to the traditional 2D system, where our system only required 1.14 tries on average while the 2D system required 2.30 tries on average. We posit that these differences came because in our system, participants were able to navigate the robot directly in the real world compared to the 2D system where participants needed to shift their attention and did the calculation based on the hint from the map.
V-C3 Subjective Evaluation of Usability
The result for the overall score and the score per question are shown in Fig. 13.
For the individual questions, statistical significance was found only in preference, professional assistance, navigation confidence, and practice necessity, suggesting that our system was generally user-friendly and facilitated easy navigation of the robot. However, as most participants lacked experience with MR head-mounted displays (HMDs), particularly the ”air tap” gesture, a learning curve was evident. This unfamiliarity with the device necessitated a period of adaptation before users could comfortably utilize the system, thus resulting the higher score for professional assistance and practice necessity. For the remaining questions, the variability in how quickly individuals learned the ”air tap” gesture introduced a subjective element to the responses, potentially contributing to the lack of statistical significance observed in these cases which contributed further to no statistical significance for the overall score.
@@ VI Conclusion
This paper has proposed MRNaB, an interface for navigating robot by using MR-beacon with Hololens 2. The establishment is done by using ”air tap” gesture with 4 functionalities which are ”add”, ”move”, ”select”, and ”delete”. MRNaB is also compacted with the database system, which removes the need for users to make the new MR-beacon when they run the project again. Based on our experiment with the traditional 2D system, our system was proven to be more effective in navigating the robot to the destination, especially in setting the beacon and navigation number. MRNaB was also proven to increase user’s confidence in order to navigate the robot to the right direction.

### v2

@@ V-A Experiment Setup
Fig. 12 shows the experiment environment. We made the environment to be similar to the household setting. For the task, we asked the participant to navigate the robot using beacon to a predetermined area with the correct direction. This area was divided into 4 stages, which were stage 1, 2, 3, and 4.
In stage 1, we wanted to cover a reasonably spacious area, that resembled a kid’s area where there is usually no clue (any reflection of object near the allowed area) from the map created by SLAM. In stage 2, we reduced the size of the area, however, we gave the opportunity for the participant to be able to see the clue which was the cabinet reflected from the map created by SLAM. This stage resembled the coffee station in the household environment. In stage 3, we kept the similar size of the area, while starting to reduce the clue reflected from the map created by SLAM, which was just a small reflection of the table . This stage resembled the working seat. Lastly in stage 4, we removed any clue from the area surrounding the target area, thus making the participant needed to use the information from the previous stage to create the MR-beacon in 2D system. This stage resembled the case where the robot needs to get the water from the water leakage from the roof.
Each of these stages was marked by green tape, and participants had to navigate the robot within the designated area. If the robot stopped or deviated outside the allowed area, which was defined as any area beyond the outer edge of the green tape, participants had to navigate the robot again using the beacon.
Fig. 13 shows the experiment flow. Each of the experiment was conducted as follows: First, participants would be given an explanatory video on how to use each of the systems. After that, we helped the participants to get used to each of the main functions. Next, we let the participants to try the system themselves. Last was the real experiment. Users are required to create a beacon by Add Button and then navigate robot to each stages by using Select Button. Every time the robot was navigated outside the area, participants would be required to move the beacon by using Move Button and navigate it again.
@@ V-B Evaluation Indices
Evaluation indices would be divided into 2, which were objective indices and subjective indices. For objective indices, we would first measure the total action number before navigation measurement which refers to the total number of clicks on the Add and Move buttons. This was the total action needed before participants finally be able to navigate the robot to the correct place. Second, we would measure navigation number measurement results per task which was the total navigation done by user per stage. This reflects how many times the user needed to adjust the robot’s placement before completing the task. Third, we would measure the action time measurement result. The total action time refers to the total time needed during Add and Move mode.
TABLE II: SUS Questionnaire Question List
No. Question
1 I think that I would like to use this system frequently.
2 I found the system unnecessarily complex.
3 I thought the system was easy to use.
4 I think that I would need the support of a technical person to be able
to use this system.
5 I found the various functions in this system were well integrated.
6 I thought there was too much inconsistency in this system.
7 I would imagine that most people would learn to use this system very
quickly.
8 I found the system very cumbersome to use.
9 I felt very confident using the system to navigate the robot.
10 I needed to learn a lot of things before I could get going with this system.
For subjective indices, we measured usability using the System Usability Scale (SUS) questionnaire [36]. The questionnaire consisted of 10 questions related to system usability, as listed in Table II. Odd-numbered questions were phrased in a positive context, while even-numbered questions were phrased in a negative context. Each question was rated on a scale from 0 to 4, and the total score was then multiplied by 2.5, resulting in a final score ranging from 0 to 100.
V-C Experiment Result
V-C1 Total Action Number Before Navigation Measurement Result
The result is shown in Fig. 14. Given that the $p$-values for stages 1, 2, 3, and Overall were below 0.05, we concluded there was statistical significance across these stages. Referencing Table III, it was evident that participants required less beacon actions to navigate the robot accurately using our system compared to a traditional 2D system where our system only required 1.59 tries on average while the 2D system required 2.95 tries on average. We posit that these changes came because, in our system, participants were able to navigate the robot directly in the real world compared to the 2D system thus making it easier for participants to navigate the robot directly to the desired location.
@@ V-C1 Total Action Number Before Navigation Measurement Result
TABLE III: Effectiveness Comparison by Action Number
@@
TABLE IV: Effectiveness Comparison by Navigation Number
@@
Figure 16: Action Time Measurement Result. The statistical difference can be found in stages 1, 2, 3, 4, and overall.
TABLE V: Effectiveness Comparison by Action Time
Stage Number Baseline (2D) MRNaB
1 9.13 14.36
2 9.84 18.79
3 8.37 13.38
4 8.87 14.46
Overall 9.05 15.25
@@ V-C2 Navigation Number Measurement Results Per Task
The result is shown in Fig. 15. Given that the $p$-values for all stages were lower than 0.05, we concluded that there is statistical significance in the data presented. Referencing Table IV, it was evident that participants required less navigation for the robot to reach its destination using our system compared to the traditional 2D system, where our system only required 1.14 tries on average while the 2D system required 2.30 tries on average. We posit that these differences came because in our system, participants were able to navigate the robot directly in the real world compared to the 2D system where participants needed to shift their attention and did the calculation based on the clue from the map.
V-C3 Action Time Measurement Result
The result is shown in Fig. 16. Given that the $p$-values for all stages were lower than 0.05, we concluded that there is statistical significance in the data presented. Referencing Table V, it was evident that participants required less action time to complete the task using traditional 2D system compared to our system. Our system only required 15.25 seconds action time on average, while the 2D system required 9.05 seconds action time on average. We posit that these differences came because in our system participants could see the targeted location in the real world, leading them to take more time to place it accurately. In contrast, the traditional system did not provide a visible target location, causing participants to place the beacon more quickly.
However, despite these drawbacks, as observed in Section V-C1 and Section V-C2, users performed fewer beacon-related actions and navigation attempts, suggesting improved efficiency in directing the robot to its destination. This indicates that while our system required more time per action, it ultimately reduced redundant interactions, enabling users to achieve their goals with fewer overall steps.
V-C4 Subjective Evaluation of Usability
The result for the overall score and the score per question are shown in Fig. 17. For the individual questions, statistical significance was found only in preference, professional assistance, navigation confidence, and practice necessity, suggesting that our system was generally user-friendly and facilitated easy navigation of the robot. We posit that the ability to see the beacon throughout the entire process increased the confidence and preference of users. Additionally, we posit that viewing detailed environmental information, rather than the abstract representation provided by a 2D map, led to higher scores for these indices.
However, as most participants lacked experience with MR head-mounted displays (HMDs), particularly the “air tap” gesture, a learning curve was evident. This unfamiliarity with the device necessitated a period of adaptation before users could comfortably utilize the system, thus resulting the higher score for professional assistance and practice necessity. For the remaining questions, the variability in how quickly individuals learned the “air tap” gesture introduced a subjective element to the responses, potentially contributing to the lack of statistical significance observed in these cases which contributed further to no statistical significance for the overall score.
@@ VI Conclusion
This paper has proposed MRNaB, an interface for navigating robot by using MR-beacon with Hololens 2. The establishment is done by using “air tap” gesture with 4 functionalities which are “Add”, “Move”, “Select”, and “Delete”. MRNaB is also integrated with a database system, eliminating the need for users to create a new MR-beacon each time the project is restarted. This ensures multi-destination visualization and maintains data persistence across sessions. Our experiments comparing MRNaB with a traditional 2D system demonstrated that despite users taking longer time to place the beacon accurately, MRNaB required less beacon settings frequency which significantly decreased the amount of navigation needed. MRNaB was also proven to increase user’s confidence in order to navigate the robot to the right direction.
There are still a few limitations to the current work. For instance, occlusion can lead to optical illusions that may mislead users when utilizing our proposed method. Additionally, establishing the MR-beacon at very far distance is challenging because the decreasing angle between the viewpoint and the floor makes it difficult to specify the location. We plan to address these issues in future work by implementing solutions for occlusion and incorporating a minimap for far-distance establishment.